OpenStudy (anonymous):

What is the rate of change for f(x) = −2 cos(4x − 3) on the interval from x = pi / 4 to x = pi / 2? A. -pi / 16 B. -16 / pi C. 2 / pi D. pi / 2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@desuoverlord

OpenStudy (anonymous):

@precal

OpenStudy (precal):

just sub and use the slope formula

OpenStudy (anonymous):

I have no idea on how to sub

OpenStudy (precal):

really this is an advance math problem......precal or calculus level

OpenStudy (precal):

strange, maybe you are in the wrong course

OpenStudy (anonymous):

im taking an algebra 2 honors class

OpenStudy (precal):

still you have sub before example f(x)=2x+3 find f(4)

OpenStudy (anonymous):

11

OpenStudy (precal):

see you know how to sub

OpenStudy (precal):

sub pi/4 into f(x)

OpenStudy (anonymous):

4.57079...

OpenStudy (precal):

no leave in radian

OpenStudy (anonymous):

(pi / 2) + 3

OpenStudy (anonymous):

so D!

OpenStudy (precal):

did you do the second substitution?

OpenStudy (precal):

pi over 4?

OpenStudy (anonymous):

what? I got pi / 2

OpenStudy (precal):

no remember you need to use the slope formula with both of them

OpenStudy (anonymous):

so I still need to find 'y'.

OpenStudy (precal):

[f(π/2)−f(π/4)]/[(π/2)−(π/4) ]

OpenStudy (precal):

remember difference of y over difference of x, slope formula

OpenStudy (anonymous):

1

OpenStudy (anonymous):

is that right??

OpenStudy (precal):

no, I don't think so, is it cos(4x-3) or cos(4x)-3 ?

OpenStudy (anonymous):

-2cos(4x - 3)

OpenStudy (precal):

need to check with my graphing calculator

OpenStudy (precal):

none are working out, are you sure you wrote the problem correctly?

OpenStudy (precal):

@bahrom7893 can you double check me?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

What is the rate of change for f(x) = −2 cos (4x − 3) on the interval from x = pi over 4 to x = pi over 2?

OpenStudy (bahrom7893):

(f(x2) - f(x1))/(x2 - x1) should give you the answer

OpenStudy (anonymous):

that makes no sense, are you solving for x?

OpenStudy (anonymous):

This is also a strangely worded question, do you mean AVERAGE rate of change? The rate of change over that interval is defined using the derivative and is used to show how the function is changing at a specific point. The average rate of change is 4cos(3)~-4, which is what these guys already said.

OpenStudy (precal):

when I did the pi/2 sub I got -2cos(4(pi/2)-3) -2 cos(2pi-3) can I simplify it anymore?

OpenStudy (precal):

but she said she is in Algebra II honors, that doesn't make any sense

OpenStudy (precal):

Algebra II doesn't do derivatives

OpenStudy (anonymous):

yeah, I didn't use it, just confirming what the question was, usually you would say average rate of change for this type of question.

OpenStudy (anonymous):

this is the final chapter in the class which is all about trig

OpenStudy (anonymous):

^^so how would you figure out to answer

OpenStudy (precal):

are you sure your problem is not -2cos(4x)-3 because it makes a big difference

OpenStudy (anonymous):

the orig. problem has no parenthesis

OpenStudy (anonymous):

http://i.imgur.com/k5T5d62.png I have done it via calculator, there is no way it is one of those answers.

OpenStudy (precal):

I believe it is what I wrote -2cos(4x)-3 -2cos(4(pi/2))-3 -2 cos2pi -3 -2(0)-3 0-3 -3

OpenStudy (anonymous):

could it be B, since there is a neagative , 16 on top, and pi on the bottom?

OpenStudy (precal):

-2cos(4x)-3 -2cos(4(pi/4))-3 -2 cos(pi )-3 -2(-1)-3 2-3 -1 we don't guess in math

OpenStudy (precal):

-3--1=-3+1=-2 on top

OpenStudy (precal):

(pi/2)-(pi/4) 2pi/4 - pi/4 pi/4

OpenStudy (precal):

-2 over pi/4 -2 times 4/pi -8/pi

OpenStudy (precal):

strange none look like that

OpenStudy (anonymous):

x2 = B

OpenStudy (precal):

@sylbot I got what you got earlier with my graphing calculator, see any mistakes on my part

OpenStudy (anonymous):

There must be an error in precal's work, I'm getting B as the answer, adjusting for the parenthesis (the -3 is outside) let me work it out tho. http://i.imgur.com/Tjrb6bK.png Ignore the bottom.

OpenStudy (precal):

yes pi over 2 is -5

OpenStudy (precal):

-4 over pi/4 -4 times 4/pi -16/pi

OpenStudy (anonymous):

thanks so much! do youhave time for one more question

OpenStudy (precal):

yes it is B btw the -3 is outside of the ( ) and that does matter

OpenStudy (anonymous):

I found the error, precal cos(pi)=1 not zero.

OpenStudy (precal):

correct its -1 and I found the mistake as well close this problem and start a new thread

OpenStudy (anonymous):

oo0ooooops cos(2pi)=1

OpenStudy (anonymous):

believe it is what I wrote -2cos(4x)-3 -2cos(4(pi/2))-3 -2 cos2pi -3 -2(0)-3 0-3 -3 here it was haha could happen to anyone, glad you got the help you need in the end.

OpenStudy (precal):

@sylbot cos (2 pi) is 1 -2(1)-3 -2-3 -5 and this is where I made the mistake but you were right thanks for catching it